OpenStudy (anonymous):

check my answer please. 3x-1/4 - x-5/5>-2

OpenStudy (anonymous):

i got x<-5

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{3x-1}{4} -\frac{x-5}{5}>-2 }\) is this what your equation is like?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{3x-1}{4} -\frac{x-5}{5}>-2 }\) common denominator is 20, so the first fraction must be multiplied times 5 on top and bottom, and second fraction must be multiplied times 4 on top and bottom. This multiplication will give you: \(\large\color{black}{ \displaystyle \frac{15x-5}{20} -\frac{4x-20}{20}>-2 }\) \(\large\color{black}{ \displaystyle \frac{15x-5 -(4x-20)}{20}>-2 }\) \(\large\color{black}{ \displaystyle \frac{15x-5 -4x+20}{20}>-2 }\) \(\large\color{black}{ \displaystyle 15x-5 -4x+20>-40 }\) \(\large\color{black}{ \displaystyle 15x-5 -4x>-60 }\) \(\large\color{black}{ \displaystyle 15x -4x>-55 }\) \(\large\color{black}{ \displaystyle 11x>-55 }\) \(\large\color{black}{ \displaystyle x>-5 }\)

OpenStudy (solomonzelman):

I don't know why you got x is LESS than -5, but the 5 on the left side is correct tho'.

OpenStudy (anonymous):

i understand where i went wrong, thats very helpfull

OpenStudy (solomonzelman):

That is very nice to hear :)